Morrow Corporation had only one job in process during May—Job X32Z—and had no finished goods inventory on May 1. Job X32Z was started in April and finished during May. Data concerning that job appear belowMorrow Corporation had only one job in process during May-Job X322—and had no finished goods inventory on May 1. Job X32Z was
Answer #1
Job X32Z
Beginning balance 5700
Direct material 9000
Direct labor 2700
Overhead 4700
Total 22100
Unit 170
Unit cost 130

a) Cost of goods sold = 75*130-370 = 9380

b) Finished goods = (170-75)*130 = 12350

c) Work in process = $0